factor the quadratic expression completely, and find the roots of the expression.
135x^2 - 222x + 91

Answer by checkley71(8403)

135X~2-222X+91=0 OR (15X-13)(9X-7)=0 OR 15X-13=0 OR 15X=13 OR X=13/15 &
9X-7=0 OR 9X=7 OR X=7/9